OpenStudy (anonymous):

If 9 is subtracted from two-thirds of a number, the result is 7. Find the number. the number is 3 the number is 10.67 the number is 24

OpenStudy (anonymous):

So, do you know how to start this problem? You need to put your word problem into an equation...

OpenStudy (anonymous):

9-2/3x=7

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Close. That's subtracting two thirds of a number from nine. You wanted to subtract nine from two thirds of a number.

OpenStudy (anonymous):

Close, but you're subtracting 9 *from* 2/3 of a number... Whoops, @nbouscal beat me to it ;)

OpenStudy (anonymous):

Let x represent the unknown number. 2/3 X - 9 = 7 2/3 X - 9 + 9 = 7 + 9 2/3 X = 16 2/3 X (3/2) = 16 (3/2) X = 24 so the answer is the number is 24.

OpenStudy (anonymous):

hope this help ☺

OpenStudy (anonymous):

thanks for the help everyone